OpenStudy (paystub-2):

2m-29-8m=-35-28m-10, anybody plz?

OpenStudy (anonymous):

solve for m?

OpenStudy (anonymous):

first combine the like terms on each side of the equation

OpenStudy (anonymous):

then move constant terms to one side and your variable terms to the other side.

OpenStudy (paystub-2):

so 2m - 8m - 10m?

OpenStudy (anonymous):

well.. actually you would combine 2m - 8m on the one side.. and -35 - 10 on the other side.

OpenStudy (anonymous):

so it would look like -29 - 6m = -45 -28m

OpenStudy (anonymous):

with me?

OpenStudy (paystub-2):

your rit my bad

OpenStudy (anonymous):

np, not your bad :) is why we are all here, for help. so you should get m = to a fraction of some sort. lemme know what u get or if you got it from here.

OpenStudy (paystub-2):

wouldnt 2m - 8m = -6?

OpenStudy (anonymous):

no, you cannot get rid of the m term by subtracting. if m was 5 then would 2(5) - 8(5) = 6? no that would be like negative 30.

OpenStudy (paystub-2):

wait i think i see how it is, so if you get, for xample, -6 and 5 was in front of it u wud put down 5-6 and not 5 - (-6)? rit?

OpenStudy (paystub-2):

are yhu buyin what im sellin?

OpenStudy (anonymous):

-6 +5 = 5-6 correct.. they both equal negative 1

OpenStudy (anonymous):

so for this question -29 - 6m = -45 - 28m now, add 28m to both sides.. u get -29 + 22m = -45. now add 29 to both sides.. you get 22m = -16 right?

OpenStudy (anonymous):

you have now have m on one side and the constant on the other. now isolate m by using division.

OpenStudy (paystub-2):

so its 22m/22 = -16/22 ?

OpenStudy (anonymous):

yep, then reduce to m = -8/11

OpenStudy (anonymous):

make sense?

OpenStudy (paystub-2):

wait how did yhu get -8/11?

OpenStudy (paystub-2):

hellooo...?

OpenStudy (anonymous):

cause 22m / 22 = -16/22 right?

OpenStudy (anonymous):

so \[m=\frac{ -16 }{ 22 }\] and if you reduce.. divide by 2 to numerator and denominator.. you get \[\frac{ -8 }{ 11 }\]

OpenStudy (paystub-2):

yes dats true thanks

OpenStudy (anonymous):

great, hope it helped. let me know if you get stuck on another :)

OpenStudy (paystub-2):

but i dont need it simplifyed i need help to get to the answer to m what does -16/22 cuz i get a wierd answer.

OpenStudy (paystub-2):

helloooooo..

OpenStudy (anonymous):

wait.. what? you need -16/22 in decimal form?

OpenStudy (anonymous):

\[-\frac{ 16 }{ 22 }=-16\div22=-.72 repeating\]